OpenStudy (lifeisadangerousgame):

Thank you FoolForMath, 14yamaka, Satelite73, Sasogeek, Asnaseer, Lukecrayonz, Mathhasproblems, Ontour, Mattt9, Magepkr728, Mertsj, UnkleRhaukus, and Zed for helping me get a 100% on my math test!! (And just so this is still technically a question) What's 3+9+300494+12+235 X 0?

OpenStudy (unklerhaukus):

3+9+300494+12+235 X 0 =12+300506+0 =300518

OpenStudy (lifeisadangerousgame):

:P actually it'd be 0 since it's all those random numbers times 0 :D

OpenStudy (unklerhaukus):

well you didn't include the bracket

OpenStudy (unklerhaukus):

(3+9+300494+12+235) X 0 = 0 but that is not the same as what yo asked

OpenStudy (lifeisadangerousgame):

oh true :/